# AlphaZero Neural Scaling and Zipf’s Law: a Tale of Board Games and Power Laws

### v1

@@ Abstract
Neural scaling laws are observed in a range of domains, to date with no clear understanding of why they occur. Recent theories suggest that loss power laws arise from Zipf’s law, a power law observed in domains like natural language. One theory suggests that language scaling laws emerge when Zipf-distributed task quanta are learned in descending order of frequency. In this paper we examine power-law scaling in AlphaZero, a reinforcement learning algorithm, using a theory of language-model scaling. We find that game states in training and inference data scale with Zipf’s law, which is known to arise from the tree structure of the environment, and examine the correlation between scaling-law and Zipf’s-law exponents. In agreement with quanta scaling theory, we find that agents optimize state loss in descending order of frequency, even though this order scales inversely with modelling complexity. We also find that inverse scaling, the failure of models to improve with size, is correlated with unusual Zipf curves where end-game states are among the most frequent states. We show evidence that larger models shift their focus to these less-important states, sacrificing their understanding of important early-game states.
@@ 1 Introduction
Neural scaling laws, describing the scaling of model performance with training resources, have been documented across many architectures and use cases (Kaplan et al., 2020; Hoffmann et al., 2022; Henighan et al., 2020; Zhai et al., 2022; Neumann & Gros, 2022). The performance of these models, typically measured by test loss, follows a power law in either compute, model size or training data. The robustness of these power laws and their usefulness in training large models, especially large language models (LLMs) (Wei et al., 2022), prompted a range of attempts to find a model explaining the mechanism behind neural power-law scaling (Sharma & Kaplan, 2020; Bahri et al., 2021; Hutter, 2021; Song et al., 2024). Several models connect performance power laws to Zipf’s law, a universal power law appearing in natural language (Michaud et al., 2023; Hutter, 2021; Bordelon et al., 2020; Cui et al., 2021; Maloney et al., 2022; Cabannes et al., 2023).
Scaling laws in reinforcement learning (RL) are so far relatively rare compared to supervised learning (Hilton et al., 2023; Tuyls et al., 2023; Springenberg et al., 2024). AlphaZero, a multi-agent RL algorithm that beat human champions in games like chess and Go (Silver et al., 2017a), exhibits power laws of Elo score with training resources, as well as a power law relation between compute and optimal model size (Neumann & Gros, 2022; Jones, 2021; Neumann & Gros, 2021).
In this paper, we explore the mechanism behind RL scaling laws, often comparing it to the quantization model of LLM scaling (Michaud et al., 2023). Our main findings are:
@@
    AlphaZero agents that exhibit scaling laws (Neumann & Gros, 2022) produce train and test data that follows Zipf’s law, suggesting learned tasks scale in a similar way. We show how agent strategies smooth and augment the Zipf power law that is known to originate from the tree-structure of board games.
@@
    The correlation between the Zipf curve and scaling exponents can be calculated by modulating policy temperature.
@@
    In agreement with quanta scaling theory, agents minimize loss on game states in descending order of frequency, counter-intuitively achieving better performance on harder tasks.
@@
    Inverse scaling, the abrupt failure of scaling laws at large sizes, coincides with an unusual state-frequency distribution. Large models scale negatively on games with an unusual tree structure that increases the frequency of strategically-unimportant late-game states. Larger agents achieve better loss on these late-game states, sacrificing performance on more important early-game states. A combination of the quantization model and selective non-stationarity of targets is suggested to explain inverse scaling in AlphaZero.
@@ 2 Background
Figure 1: Zipf’s law in AlphaZero games. Board-state frequency follows a power law in state rank, here for Connect Four and Pentago. Similar exponents $\alpha$ appear for agents of various sizes trained on different games, see appendix B.1.
A short description of the AlphaZero algorithm is given in appendix A.
@@
Zipf’s law is an empirical law describing the distribution of values when sorted in decreasing order, typically with respect to frequency (Zipf, 2013). When sorting elements in decreasing order of frequency, e.g. words in natural language texts, one regularly finds that the frequency distribution $S(n)$ follows a power law in element rank $n$:
@@
often with $\alpha\approx 1$ for natural language datasets (Moreno-Sánchez et al., 2016; Piantadosi, 2014).
@@
Michaud et al. (2023) propose the quantization model to explain the origin of power-law scaling in LLMs, connecting neural-network training to language Zipf’s laws. This model assumes that LLMs learn discreet task quanta, and that the frequency of these tasks follows Zipf’s law. The quantization model suggests that language models learn tasks in descending order of frequency, which leads to power-law scaling laws. If the loss on each task is reduced by a fixed amount $\Delta L$ after it is learned, then a model that learned the first $n$ tasks will have an expected loss of
$L=\frac{\Delta L}{\alpha\zeta(\alpha+1)}\cdot n^{1-\alpha}+L_{\infty}\,\implies\,(L-L_{\infty})\propto n^{1-\alpha}\,.$ (2)
Where $\zeta$ is the Riemann zeta function. It is assumed that a fixed neural capacity $c$ is needed to fit each of the independent task quanta. At the limit of infinite compute and data, the number of learned task quanta $n=c\cdot N$ is determined by the number of parameters $N$. Eq. 2 then becomes a power-law scaling law:
@@
A large body of work exists on models of neural scaling laws on static datasets, several of which making the connection between Zipf’s law and scaling laws. Hutter (2021) show how a data scaling law emerges when training on Zipf-distributed data under the assumption of infinite memory. Bordelon et al. (2020) develop a model of data scaling laws for kernels, which applies to neural nets at the infinite width limit. Other works expand this line of work on kernel models, including results on real datasets (Cui et al., 2021; 2023). Maloney et al. (2022) develop a model that explains joint scaling of model size and dataset size. Cabannes et al. (2023) assume Zipf-distributed data and develop a model for scaling of attention models. Dohmatob et al. (2024) explore model collapse through the effect of synthetic data on heavy-tailed data distributions.
We frequently compare our results to Michaud et al. (2023) throughout this paper, since their model develops the full set of scaling laws for neural nets (compute, parameters, data) by assuming a Zipf’s law of the data. Our comparisons also hold for the above-mentioned works, depending on the assumptions they make. To our knowledge, no complete model exists that explains scaling laws in any RL setting.
@@ 3 Related work
To our knowledge, no study has been made on Zipf’s law in AI games. Blasius & Tönjes (2009) and Georgeot & Giraud (2012) show that the popularity of openings in human game datasets follows Zipf’s law in chess and Go, respectively. Blasius & Tönjes (2009) are the first to propose a connection to the branching-tree structure of chess, without proof. We provide evidence supporting such a connection in section 5.1.
@@
In sections 6 and 7 we present results on the scaling of AlphaZero loss on individual states. A similar analysis is done by McGrath et al. (2022) and Lovering et al. (2022), who both analyze how human concepts are learned by AlphaZero by measuring accuracy. McGrath et al. (2022) map how human chess concepts are learned, both as a function of training length and neural network depth. They measure the accuracy of a few hand-crafted concepts, compared to our analysis that measures loss on up to $10^{5}$ states. Similar to our inverse-scaling results in section 7, Lovering et al. (2022) observe differences between how short-term and long-term concepts are learned in the game of Hex.
@@
Our work discusses the effects of state frequency in the training data on AlphaZero performance. Another observation on this relation of data frequency with performance is found in Ruoss et al. (2024), where Transformers are trained with supervised learning on annotated human chess games. The authors perform an ablation study where they change the frequencies of board states in the dataset, either sampling states uniformly or sampling them according to the frequency they were played by humans, which should follow Zipf’s law. Testing on chess puzzle accuracy, changing the state frequency seems to have a strong effect on model performance, in favor of uniform sampling.
@@ 4 Methods
We use AlphaZero agents trained on four board games: Connect Four, Pentago, Oware and Checkers. All neural nets are either taken from Neumann & Gros (2022) or trained using the same code, which is based on OpenSpiel (Lanctot et al., 2019). We keep the same hyperparameters, training MLP agents with 3 hidden layers at exponentially-increasing widths, with a fixed amount of MCTS steps per move (300). We use the open-source model weights for Connect Four and Pentago, and train new models on Oware and Checkers. All of these models exhibit power-law scaling laws, either partially or fully: Connect Four and Pentago scaling laws are discussed at length in Neumann & Gros (2022), while we present new size-scaling curves for Oware and Checkers in Fig. 5A. The power laws are in the Bradley-Terry playing strength $\gamma$, which determines the Elo score: $\text{Elo}\propto\log_{10}(\gamma)$ (Bradley & Terry, 1952; Elo, 1978). $\gamma$ scales as a power of neural-net parameters and training compute (Neumann & Gros, 2022):
@@
Most results in this paper involve the frequency of states $s\in\mathcal{S}$ in a Markov decision process $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}\rangle$ (Puterman, 2014). In practice, we identify each state by its observation tensor, which is the input to the neural net. In all four games, this observation is the position of each piece on the board, and does not include the history of previous moves (unlike the chess analysis of Blasius & Tönjes (2009)).
5 Uncovering Zipf’s law in AlphaZero
5.1 Zipf’s law in games with scaling laws
@@
We find a Zipf distribution in games played by AlphaZero agents that exhibit scaling laws. Fig. 1 shows the frequency distribution of board states played by Connect Four and Pentago AlphaZero agents of varying sizes during training. The number of times each state was visited follows a clear Zipf power law when sorted in descending order. The exact value of the exponent defined in Eq. 1 fluctuates in the range $\alpha\sim 0.8\!-\!1.0$ for different models and games, close to the value observed in human Go games (Georgeot & Giraud, 2012) and to the exponent of random games, discussed in the next section. Appendix B.1 contains plots of individual-agent Zipf curves, and discusses the variation of the power-law exponent $\alpha$. The plateaus at the tail-end of each distribution are due to finite-size effects, and smooth-out when more state-frequency data is available.
This result strengthens the assumption that Zipf’s law is not caused by human decision-making, but is rather derived from the game rules. The existence of Zipf’s law in human games has long since been established, but its cause is not entirely clear. The frequency of popular opening moves in games like Chess and Go follows a descending power law in rank, both for expert and amateur human players (Blasius & Tönjes, 2009; Georgeot & Giraud, 2012). The fact that an RL model trained with no human knowledge produces Zipf’s law similar to humans implies that Zipf’s law is not caused by human strategy preference.
@@
We give some intuition to the source of Zipf’s law in Fig. 2. The relation of Zipf’s law to the underlying branching process of board games was already pointed out by Blasius & Tönjes (2009). It is easy to show that an ideal game, with a constant branching factor and a tree structure free of loops, would produce a Zipf’s law distribution. Specifically, one can show that the frequency distribution in such a game is a series of plateaus around the line $S(n)\propto n$. The same reasoning explains the appearance of Zipf’s law in randomly-generated texts (Li, 1992). In Fig. 2A, we visualize this ideal-game Zipf’s law by plotting the state-frequency distribution for a non-looping, constant branching game with fixed length. Increasing the branching factor widens the plateaus, spreading them out. We use the initial branching factor of Connect Four.
In Fig. 2B we show that adding game rules molds the plateaus into a smoother curve. Using a uniformly random policy to play Connect Four and Pentago smooths the power-law-centered plateaus, somewhat merging them into a straight line. The toy-model exponent changes slightly after adding game rules, but the general power-law trend remains intact.
Appendix B.2 contains additional results on the emergence of Zipf’s law, and shows that non-uniform policies can also smooth the distribution into a power law. As we see in Fig. 1, a learned policy smooths the distribution further, keeping the power-law shape while slightly changing the exponent.
5.2 Temperature and correlation with scaling laws
@@
where $N(s,a)$ is the number of times the state-action pair $(s,a)$ was probed, see appendix D for a detailed explanation. The temperature $T$ interpolates between a deterministic policy choosing the best move at $T=0$, and a uniform policy at $T=\infty$. At inference time, e.g. when calculating Elo ratings, $T$ is either set to zero or reduced to a low value, the latter to avoid deterministic policies.
@@
Additional experiments confirm that the relation between the state distribution and the scaling law is causal, meaning the scaling law changes only because of the changing Zipf distribution. A possible non-causal effect can be Elo scores changing due to the degradation of the policy $\bm{\pi}$ at high $T$. We measure the effect of $T$ on policy quality in appendix C and find that the probability policy $\bm{\pi}(T)$ assigns to optimal moves only starts decreasing at $T>0.5$. The correlation in Fig. 2C is strictly causal, since it falls in the temperature range where policy quality is preserved. Interestingly, the Zipf exponent only changes in the temperature where policy quality is optimal, and is fixed for $T>0.5$.

### v2

@@ Abstract
Neural scaling laws are observed in a range of domains, to date with no universal understanding of why they occur. Recent theories suggest that loss power laws arise from Zipf’s law, a power law observed in domains like natural language. One theory suggests that language scaling laws emerge when Zipf-distributed task quanta are learned in descending order of frequency. In this paper we examine power-law scaling in AlphaZero, a reinforcement learning algorithm, using a model of language-model scaling. We find that game states in training and inference data scale with Zipf’s law, which is known to arise from the tree structure of the environment, and examine the correlation between scaling-law and Zipf’s-law exponents. In agreement with the quanta scaling model, we find that agents optimize state loss in descending order of frequency, even though this order scales inversely with modelling complexity. We also find that inverse scaling, the failure of models to improve with size, is correlated with unusual Zipf curves where end-game states are among the most frequent states. We show evidence that larger models shift their focus to these less-important states, sacrificing their understanding of important early-game states.
@@ 1 Introduction
Neural scaling laws, describing the scaling of model performance with training resources, have been documented across many architectures and use cases kaplan2020scaling ; hoffmann2022training ; henighan2020scaling ; zhai2022scaling ; neumann2022scaling . The performance of these models, typically measured by test loss, follows a power law in either compute, model size or training data. The robustness of these power laws and their usefulness in training large models, especially large language models (LLMs) wei2022emergent , prompted a range of attempts to find a model explaining the mechanism behind neural power-law scaling sharma2020neural ; bahri2021explaining ; hutter2021learning ; song2024resource ; paquette20244 . Several models connect performance power laws to Zipf’s law, a universal power law appearing in natural language michaud2023quantization ; hutter2021learning ; bordelon2020spectrum ; cui2021generalization ; maloney2022solvable ; cabannes2023scaling .
Scaling laws in reinforcement learning (RL) are so far relatively rare compared to supervised learning hilton2023scaling ; tuyls2023scaling ; springenberg2024offline . AlphaZero, a multi-agent RL algorithm that beat human champions in games like chess and Go silver2017mastering2 , exhibits power laws of Elo score with training resources, as well as a power law relation between compute and optimal model size neumann2022scaling ; jones2021scaling ; neumann2021investment .
In this paper, we explore the mechanism behind known RL scaling laws, comparing this mechanism to the quantization model of LLM scaling michaud2023quantization . Our main findings are:
@@
    We find that AlphaZero agents that exhibit scaling laws, such as those trained in Neumann & Gros (2022) neumann2022scaling , produce train and test data that follows Zipf’s law, suggesting that learned tasks also scale with Zipf’s law.
@@
    We show how these smooth Zipf laws form when agent policies are combined with the tree-structure of board games, which is known to create a Zipf-like state distribution blasius2009zipf . We find a correlation between the Zipf curve and scaling exponents, exposed by modulating policy temperature at inference.
@@
    In agreement with the quanta scaling model, we show that agents minimize loss on game states in descending order of frequency, counter-intuitively achieving better performance on harder tasks.
@@
    Inverse scaling, the abrupt failure of scaling laws at large sizes, coincides with an unusual state-frequency distribution. We show that large models scale negatively on games with an unusual tree structure that increases the frequency of strategically-unimportant late-game states. Larger agents achieve better loss on these late-game states, sacrificing performance on more important early-game states. We propose an explanation of inverse scaling in AlphaZero using a combination of the quantization model and the selective non-stationarity of training targets.
@@ 2 Background
Figure 1: Zipf’s law in AlphaZero games. Board-state frequency follows a power law in state rank, here for Connect Four and Pentago. Similar exponents $\alpha$ appear for agents of various sizes trained on different games, see appendix C.1.
A short description of the AlphaZero algorithm and AlphaZero scaling laws is given in appendix A.
@@
Zipf’s law is an empirical law describing the distribution of values when sorted in decreasing order, typically with respect to frequency zipf2013psycho . When sorting elements in decreasing order of frequency, e.g. words in natural language texts, one regularly finds that the frequency distribution $S(n)$ follows a power law in element rank $n$:
@@
often with $\alpha\approx 1$ for natural language datasets moreno2016large ; piantadosi2014zipf .
@@
Michaud et al. (2023) michaud2023quantization proposed that a ‘quantization model’ may explain the origin of power-law scaling in LLMs, connecting neural-network training to language Zipf’s laws. The model assumes that LLMs learn discrete task quanta and that the frequency of these tasks follows Zipf’s law. The quantization model suggests that language models learn tasks in descending order of frequency, which leads to power-law scaling laws. If the loss on each task is reduced by a fixed amount $\Delta L$ after it is learned, then a model that learned the first $n$ tasks will have an expected loss of
$L=\frac{\Delta L}{\alpha\zeta(\alpha+1)}\cdot n^{1-\alpha}+L_{\infty}\,\implies\,(L-L_{\infty})\propto n^{1-\alpha}\,$, (2)
where $\zeta$ is the Riemann zeta function. It is assumed that a fixed neural capacity $c$ is needed to fit each of the independent task quanta. At the limit of infinite compute and data, the number of learned task quanta $n=\frac{N}{c}$ is determined by the number of parameters $N$. Eq. 2 then becomes a power-law scaling law:
@@
A large body of work exists on models of neural scaling laws on static datasets, several of which making the connection between Zipf’s law and scaling laws. Hutter (2021) hutter2021learning argues that data scaling laws emerge when training on Zipf-distributed data under the assumption of infinite memory. Bordelon et al. (2020) bordelon2020spectrum develop a model of data scaling laws for kernels, which applies to neural nets at the infinite width limit. Other works expand this line of work on kernel models, including results on real datasets cui2021generalization ; cui2023error . Maloney et al. (2022) maloney2022solvable develop a model that explains joint scaling of model size and dataset size. Cabannes et al. (2023) cabannes2023scaling assume Zipf-distributed data and develop a model for scaling of attention models. Schaeffer et al. (2025) schaeffer2025large propose that exponential scaling on individual tasks can aggregate to a power law when the distribution of success probabilities is heavy tailed, in the case of multiple attempts per task. Dohmatob et al. (2024) dohmatob2024tale explore model collapse through the effect of synthetic data on heavy-tailed data distributions.
We frequently compare our results to Michaud et al. (2023) michaud2023quantization throughout this paper, since their model develops the full set of scaling laws for neural nets (compute, parameters, data) by assuming a Zipf’s law of the data. Our comparisons also hold for the above-mentioned works, depending on the assumptions they made. To our knowledge, no complete model exists that explains scaling laws in any RL setting.
@@ 3 Related work
To our knowledge, no study has been made on Zipf’s law in AI games. Blasius & Tönjes (2009) blasius2009zipf and Georgeot & Giraud (2012) georgeot2012game show that the popularity of openings in human game datasets follows Zipf’s law in chess and Go, respectively. Blasius & Tönjes are the first to propose a connection to the branching-tree structure of chess, without proof. We provide evidence supporting such a connection in section 5.
@@
In sections 7 and 8 we present results on the scaling of AlphaZero loss on individual states. A similar analysis is done by McGrath et al. (2022) mcgrath2022acquisition and Lovering et al. (2022) lovering2022evaluation , who both analyze how human concepts are learned by AlphaZero by measuring accuracy. McGrath et al. (2022) mcgrath2022acquisition map how human chess concepts are learned, both as a function of training length and neural network depth. They measure the accuracy of a few hand-crafted concepts, in contrast to our analysis that measures loss on up to $10^{5}$ unique states. Similar to our inverse-scaling results in section 8, Lovering et al(̇2022) lovering2022evaluation observe differences between how short-term and long-term concepts are learned in the game of Hex.
@@
Our work discusses the effects of training-data state frequency on AlphaZero agent performance. Another observation regarding the relation of data frequency with performance can be found in Ruoss et al. (2024), ruoss2024grandmaster , where Transformers are trained with supervised learning on annotated human chess games. The authors perform an ablation study where they change the frequencies of board states in the dataset, either sampling states uniformly or sampling them according to their frequency in human games, which follows Zipf’s law. Testing on chess puzzle accuracy, they find that changing the state frequency has a strong effect on model performance, favoring uniform sampling over sampling from the Zipf distribution.
@@ 4 Methods
We use AlphaZero agents trained on four board games: Connect Four, Pentago, Oware and Checkers. All neural nets are either open-source models trained in Neumann & Gros (2022) neumann2022scaling , or newly trained using the same open-source OpenSpiel code LanctotEtAl2019OpenSpiel . We use the open-source model weights for Connect Four and Pentago, and train new models with the same hyperparameters on Oware and Checkers. All of these models exhibit power-law scaling laws, either partially or fully: Connect Four and Pentago scaling laws are discussed at length at Neumann & Gros neumann2022scaling , here we present size-scaling curves for Oware and Checkers, see Fig. 5A. The power laws are for the Bradley-Terry playing strength $\gamma$, which determines the Elo score: $\text{Elo}\propto\log_{10}(\gamma)$ bradley1952rank ; elo1978rating . One finds that $\gamma$ scales as a power of neural-net parameters and training compute neumann2022scaling :
@@
Most results in this paper involve the frequency of states $s\in\mathcal{S}$ in a Markov decision process $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R}\rangle$ puterman2014markov . In practice, we identify each state by its observation tensor, which is the input to the neural net. In all four games, this observation is the position of each piece on the board, and does not include the history of previous moves (unlike the chess analysis of Blasius & Tönjes (2009) blasius2009zipf ). All AlphaZero agents use MLP nets with 3 hidden layers at exponentially-increasing widths, with a fixed amount of 300 MCTS steps per move during both training and inference.
5 Zipf’s law in AlphaZero
@@
We find a Zipf distribution in games played by AlphaZero agents that exhibit scaling laws. Fig. 1 shows the frequency distribution of board states played by Connect Four and Pentago AlphaZero agents of varying sizes during training. The number of times each state was visited follows a clear Zipf power law when sorted in descending order. The exact value of the exponent defined in Eq. 1 fluctuates in the range $\alpha\sim 0.8\!-\!1.0$ for different models and games, close to the value observed in human Go games georgeot2012game and to the exponent of random games, discussed in the next section. Appendix C.1 contains plots of individual-agent Zipf curves, and discusses the variation of the power-law exponent $\alpha$. The plateaus at the tail-end of each distribution are due to finite-size effects, and smooth-out when more state-frequency data is available.
This result strengthens the assumption that Zipf’s law is not caused by human decision-making, but is rather derived from the game rules. The existence of Zipf’s law in human games has long since been established, but its cause is not entirely clear. The frequency of popular opening moves in games like Chess and Go follows a descending power law in rank, both for expert and amateur human players blasius2009zipf ; georgeot2012game . The fact that an RL model trained with no human knowledge produces Zipf’s law similar to humans implies that Zipf’s law is not caused by human strategy preference.
@@
A first intuition regarding the source of Zipf’s law is presented in Fig. 2. The relation of Zipf’s law to the underlying branching process of board games was already pointed out by Blasius & Tönjes (2009) blasius2009zipf . It is easy to show that an ideal game, with a constant branching factor and a tree structure free of loops, would produce a Zipf-like distribution. Specifically, one can show that the frequency distribution in such a game is a series of plateaus around the line $S(n)\propto n$, see appendix J. The same reasoning explains the appearance of Zipf’s law in randomly-generated texts li1992random . In Fig. 2A this ideal-game Zipf’s law is visualized by plotting the state-frequency distribution for a non-looping, constant branching game with fixed length. Increasing the branching factor widens the plateaus, spreading them out. We use the initial branching factor of Connect Four.
In Fig. 2B we show that adding game rules molds the plateaus into a smoother curve. Using a uniformly random policy to play Connect Four and Pentago smooths the power-law-centered plateaus, roughly merging them into a straight line. The toy-model exponent changes slightly after adding game rules, but the general power-law trend remains intact.
Appendix C.2 contains additional results on the emergence of Zipf’s law, and shows that non-uniform policies can also smooth the distribution into a power law. As we see in Fig. 1, a learned policy smooths the distribution further, keeping the power-law shape while slightly changing the exponent.
6 Temperature and correlation with scaling laws
@@
where $N(s,a)$ is the number of times the state-action pair $(s,a)$ was probed, see appendix F for a detailed explanation. The temperature $T$ interpolates between a deterministic policy choosing the best move at $T=0$, and a uniform policy at $T=\infty$. At inference time, e.g. when calculating Elo ratings, $T$ is either set to zero or reduced to a low value, the latter to avoid deterministic policies.
@@
Additional experiments confirm that the relation between the state distribution and the scaling law is causal, meaning the scaling law changes only because of the changing Zipf distribution. A possible non-causal effect could be Elo scores changing due to the degradation of the policy $\bm{\pi}$ at high $T$. We measure the effect of $T$ on policy quality in appendix D and find that the probability the policy $\bm{\pi}(T)$ assigns to optimal moves only starts decreasing at $T>0.5$. The correlation in Fig. 2C is strictly causal, since it falls in the temperature range where policy quality is preserved. Interestingly, the Zipf exponent only changes in the temperature where policy quality is optimal, and is fixed for $T>0.5$.
@@
Temperature can augment the Zipf curve since it affects exploration, which in turn affects the state distribution, as we also show in appendix C.2. This implies that changing the MCTS parameters could also have an effect on state distribution, but we did not verify this relation.